# Steering Knowledge Selection Behaviours in LLMs via SAE-Based Representation Engineering

### v1

@@ 1 Introduction
Large language models (LLMs) have shown remarkable capability to memorise factual knowledge and solve knowledge-intensive tasks (Petroni et al., 2019; Brown, 2020; Touvron et al., 2023; Jiang et al., 2023; Anil et al., 2023). Nevertheless, the knowledge stored in their parameters (parametric knowledge) can be inaccurate or outdated (Xu et al., 2024). To alleviate this issue, retrieval and tool-augmented approaches have been widely adopted to provide LLMs with external knowledge (contextual knowledge(Karpukhin et al., 2020; Lewis et al., 2020; Wu et al., 2022; Schick et al., 2024). However, contextual knowledge may sometimes conflict with the parametric knowledge of the model, leading to what we refer to as knowledge conflicts. Such conflicts can cause undesired behaviour, where the model may rely on inaccurate information sources, resulting in incorrect outputs (Mallen et al., 2023; Xie et al., 2024a; Su et al., 2024; Wang et al., 2023; Zhao et al., 2024).
@@ 3 Detection of Knowledge Conflicts
We apply a linear probing method (Conneau et al., 2018; Zhu and Li, 2023; Allen-Zhu and Li, 2023) to investigate whether the residual stream contains a signal of knowledge conflict. Specifically, we train logistic regression models to classify whether a given activation (the hidden state, MLP or Self-Attention activations) is from the $\mathcal{D}_{E_{C}}$ or $\mathcal{D}_{E_{M}}$, i.e. whether it contains a knowledge conflict or not. We use activations from each layer as input and formulate this as a binary classification task. The evaluation is conducted on a held-out test set. We present probing results on Llama2-7B (Touvron et al., 2023) and Gemma2-9B (Rivière et al., 2024) using AUROC as metric in Fig. 2. We observe that the probing accuracy increases from the first layer to the middle layers, and this trend is the same across different types of activations. This indicates that we can detect the signal of knowledge conflict in the residual stream of the mid-layers. The probing accuracy decreases in the later layers, especially for MLP and Self-Attention activations, which indicates that MLP and Self-Attention modules do not further add the signal of conflicting knowledge to the residual stream. We provide more details and analysis about knowledge conflict detection in Appendix A.
@@ 6.2 Analysing the Residual Stream
In Fig. 5(c), we present the probing results on the residual stream using the same probing model described in Section 3. We observe that when we steer the usage of parametric knowledge, the probing performance decreases immediately (green line), indicating that the signal of knowledge conflict fades quickly, and the representations of activations become closer to $\mathcal{D}_{E_{M}}$, thus making it more difficult for the probing model to distinguish whether a given activation is from $\mathcal{D}_{E_{C}}$ or $\mathcal{D}_{E_{M}}$. In contrast, when we steer the usage of contextual knowledge, the probing performance increases (blue line), indicating the signal of the conflict increases, and the representations of activations become more different from $\mathcal{D}_{E_{M}}$, making it easier for the probing model to distinguish whether a given activations is from $\mathcal{D}_{E_{C}}$ or $\mathcal{D}_{E_{M}}$.
@@
We provide more analysis of the representation patterns in Appendix F. In this work, we only provide our empirical observation on the representation pattern and leave investigating the reasons in future works.
@@ 7 Related Works
Knowledge conflicts refer to discrepancies among contextual and parametric knowledge (Chen et al., 2022; Xie et al., 2024b). Xu et al. (2024) identify three types of knowledge conflicts: inter-context (Zhang and Choi, 2021; Du et al., 2022; Pan et al., 2023b), context-memory (Longpre et al., 2021; Xie et al., 2024b), and intra-memory conflicts (Huang et al., 2023). In this work, we focus on context-memory knowledge conflicts, which refers to conflicts between the contextual knowledge and the parametric knowledge encoded in the model parameters.
@@
Sparse Auto-Encoders (SAEs) have been introduced as a post-hoc analysis tool to identify disentangled features within uncompressed representations of an LLM (Yun et al., 2021; Bricken et al., 2023; Huben et al., 2024). SAEs are trained with sparsity regularisation to learn a sparse, overcomplete basis that characterises the activation space of an LLM (Bereska and Gavves, 2024). Marks et al. (2024) showed that the features learned by SAEs can identify sparse circuits in LLMs. Templeton et al. (2024b) showed the possibility of searching for monosemantic features and steering LLMs’ generation.
@@ 8 Conclusions
We investigated the context-memory knowledge conflicts in LLMs. We identify that knowledge conflicts can be detected by probing the residual stream of the model (Section 3) and propose SpARE (Section 4), a training-free representation engineering method that leverages pre-trained SAEs to effectively and efficiently control the knowledge selection behaviour of LLMs at inference time. Our experimental results on ODQA tasks show that SpARE produces more accurate results than existing representation engineering and contrastive decoding methods. By providing a mechanism to steer knowledge selection behaviours at inference time, SpARE offers a promising approach to managing knowledge conflicts in LLMs without significant computational overhead.
@@ A.2 More Probing Results
We present the knowledge conflict probing results of Llama2-7B on NQSwap and Macnoise on Fig. 8 and Fig. 9 using accuracy, AUROC and AUPRC as metrics.
@@ Appendix B Sparse Auto-Encoders Details
Following Gao et al. (2024), we pre-train the SAEs for Llama2-7B with TopK activation functions: $\mathbf{z}=\text{TopK}\left(\mathbf{W}_{\theta}(\mathbf{h}-\mathbf{b})\right)$, which only keeps the $k$ largest latents during pre-training, while does not use sparsity regularisation during pre-training. The loss is calculated by the L2 norm of reconstruction error: $\mathcal{L}=\|\mathbf{h}-\hat{\mathbf{h}}\|_{2}^{2}$. We pre-train SAE models [^3] https://github.com/EleutherAI/sae with $10$B tokens sampled from RedPajama [^4] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-V2 and use the hyperparamters determined by Gao et al. (2024). The pre-training for an SAE for a certain layer of hidden states costs about $300$ 80G A100 GPU hours.
@@ C.2 Identifying Functional Activations
Since a higher $I(Z_{i};Y)$555 We estimate mutual information between continuous variables $Z_{i}$ and discrete labels $Y$ using https://scikit-learn.org/1.5/modules/generated/sklearn.feature_selection.mutual_info_classif.html indicates a higher dependency between $Z_{i}$ and the knowledge selection behaviour, we sorted $\{I(Z_{i};Y)\}_{i=1}^{n}$ in descending order, and consider the top $k$ activation work as functional activations. To decide the number of selected activations $k$, we introduce a hyperparameter $K$. We then choose $k$ such that:
@@ C.3 Impact of the Size of the Collected Activations
(a) Use different numbers of activations to calculate mutual information.
(b) Use different numbers of activations to calculate $\mathbf{z}_{C}$ and $\mathbf{z}_{M}$
@@ Appendix F Distribution Patterns of the Residual Stream Under Knowledge Conflict
In this section, we provide further analysis of the representation patterns when knowledge conflicts in addition to Fig. 5(d). Here, we focus on analysing the distribution patterns of different knowledge selection behaviours. More specifically, we compare the representation difference between the activation from $\mathcal{D}_{C}$ and $\mathcal{D}_{M}$, which are both under knowledge conflict but select contextual and parametric knowledge to generate the answer $C$ and $M$. In Section F.1, we analyse the skewness patterns; in Section F.2, we analyse the L1 norm and L2 norm patterns since previous work (Devoto et al., 2024) also show the norm value may be related to the contextual information usage.

### v3

@@ 1 Introduction
Large language models (LLMs) have shown remarkable capability to memorise factual knowledge and solve knowledge-intensive tasks (Petroni et al., 2019; Brown, 2020; Touvron et al., 2023; Jiang et al., 2023; Anil et al., 2023). Nevertheless, the knowledge stored in their parameters (parametric knowledge) can be inaccurate or outdated (Xu et al., 2024). To alleviate this issue, retrieval and tool-augmented approaches have been widely adopted to provide LLMs with external knowledge (contextual knowledge(Karpukhin et al., 2020; Lewis et al., 2020; Wu et al., 2022; Schick et al., 2024). However, contextual knowledge may sometimes conflict with the parametric knowledge of the model, leading to what we refer to as knowledge conflicts. Such conflicts can cause undesired behaviour, where the model may rely on inaccurate information sources, resulting in incorrect outputs (Mallen et al., 2023; Xie et al., 2024a; Su et al., 2024; Wang et al., 2023; Zhao et al., 2024a).
@@ 3 Detection of Knowledge Conflicts
We apply a linear probing method (Conneau et al., 2018; Zhu and Li, 2023; Allen-Zhu and Li, 2023) to investigate whether the residual stream contains a signal of knowledge conflict. Specifically, we train logistic regression models to classify whether a given activation (the hidden state, MLP or Self-Attention activations) is from the $\mathcal{D}_{E_{C}}$ or $\mathcal{D}_{E_{M}}$, i.e. whether it contains a knowledge conflict or not. We use activations from each layer as input and formulate this as a binary classification task. The evaluation is conducted on a held-out test set. We present probing results on Llama2-7B (Touvron et al., 2023) and Gemma2-9B (Rivière et al., 2024) using AUROC as metric in Fig. 2. We observe that the probing accuracy increases from the first layer to the middle layers, and this trend is the same across different types of activations. This indicates that we can detect the signal of knowledge conflict in the residual stream of the mid-layers. The probing accuracy decreases in the later layers, especially for MLP and Self-Attention activations, which indicates that MLP and Self-Attention modules do not further add the signal of conflicting knowledge to the residual stream. We provide more details and analysis about knowledge conflict detection in Appendix A and Zhao et al. (2024b).
@@ 6.2 Analysing the Residual Stream
In Fig. 5(c), we present the probing results on the residual stream using the same probing model described in Section 3. We observe that when we steer towards the usage of parametric knowledge, the probing performance decreases immediately (green line), indicating that the signal of knowledge conflict fades quickly, and the representations of activations become closer to $\mathcal{D}_{E_{M}}$, thus making it more difficult for the probing model to distinguish whether a given activation is from $\mathcal{D}_{E_{C}}$ or $\mathcal{D}_{E_{M}}$. In contrast, when we steer towards using contextual knowledge, the probing performance increases (blue line), indicating the signal of the conflict increases, and the representations of activations become more different from $\mathcal{D}_{E_{M}}$, making it easier for the probing model to distinguish whether a given activations is from $\mathcal{D}_{E_{C}}$ or $\mathcal{D}_{E_{M}}$.
@@
We provide more analysis of the representation patterns in Appendix F and Zhao et al. (2024b). In this work, we only provide our empirical observation on the representation pattern and leave investigating the reasons in future works.
@@ 7 Related Works
Knowledge conflicts refer to discrepancies among contextual and parametric knowledge (Chen et al., 2022; Xie et al., 2024b). Xu et al. (2024) identify three types of knowledge conflicts: inter-context (Zhang and Choi, 2021; Du et al., 2022; Pan et al., 2023b; Zhao et al., 2024c), context-memory (Longpre et al., 2021; Xie et al., 2024b; Minder et al., 2024), and intra-memory conflicts (Huang et al., 2023). In this work, we focus on context-memory knowledge conflict, which refers to conflicts between the contextual knowledge and the parametric knowledge encoded in the model parameters. Ortu et al. (2024) and Jin et al. (2024) investigate the mechanisms of attention heads and feed-forward networks of LLMs when context-memory knowledge conflict occurs.
@@
Sparse Auto-Encoders (SAEs) have been introduced as a post-hoc analysis tool to identify disentangled features within uncompressed representations of an LLM (Yun et al., 2021; Bricken et al., 2023; Huben et al., 2024). SAEs are trained with sparsity regularisation to learn a sparse, overcomplete basis that characterises the activation space of an LLM (Bereska and Gavves, 2024). Marks et al. (2024) showed that the features learned by SAEs can identify sparse circuits in LLMs. Templeton et al. (2024b) showed the possibility of searching for monosemantic features and steering LLMs’ generation. Chalnev et al. (2024) improves steering vectors using SAEs. Gur-Arieh et al. (2025) generates description for representations learned through SEAs. (Lan et al., 2024) finds universal representation across different language models.
@@ 8 Conclusions
We investigated the context-memory knowledge conflicts in LLMs. We identify that knowledge conflicts can be detected by probing the residual stream of the model (Section 3) and propose SpARE (Section 4), a training-free representation engineering method that leverages pre-trained SAEs to effectively and efficiently control the knowledge selection behaviour of LLMs at inference time. Our experimental results on ODQA tasks show that SpARE produces more accurate results than existing representation engineering and contrastive decoding methods (Section 5). By providing a mechanism to steer knowledge selection behaviours at inference time, SpARE offers a promising approach to managing knowledge conflicts in LLMs without significant computational overhead. Additionally, we investigate the residual stream of LLMs under knowledge conflicts (Section 6). We find that 1) the knowledge selection behaviour is more steerable at the middle layers of LLMs; 2) the residual stream shows a significantly more skewed representation when models using contextual knowledge compared to using parametric knowledge.
@@ A.2 More Probing Results
We present the knowledge conflict probing results of Llama2-7B on NQSwap and Macnoise on Fig. 8 and Fig. 9 using accuracy, AUROC and AUPRC as metrics. We provide more analysis of probing the residual stream under knowledge conflict in our preliminary study (Zhao et al., 2024b).
@@ Appendix B Sparse Auto-Encoders Details
Following Gao et al. (2024), we pre-train the SAEs for Llama2-7B with TopK activation functions: $\mathbf{z}=\text{TopK}\left(\mathbf{W}_{\theta}(\mathbf{h}-\mathbf{b})\right)$, which only keeps the $k$ largest latents during pre-training, while does not use sparsity regularisation during pre-training. The loss is calculated by the L2 norm of reconstruction error: $\mathcal{L}=\|\mathbf{h}-\hat{\mathbf{h}}\|_{2}^{2}$. We pre-train SAE models [^3] https://github.com/EleutherAI/sae with $10$B tokens sampled from RedPajama [^4] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-V2 and use the hyperparamters determined by Gao et al. (2024). The pre-training for an SAE for a certain layer of hidden states costs about $300$ 80G A100 GPU hours. [^5] We provide our pre-trained SAEs in https://huggingface.co/yuzhaouoe/Llama2-7b-SAE/tree/main
@@ C.2 Identifying Functional Activations
Since a higher $I(Z_{i};Y)$666 We estimate mutual information between continuous variables $Z_{i}$ and discrete labels $Y$ using https://scikit-learn.org/1.5/modules/generated/sklearn.feature_selection.mutual_info_classif.html indicates a higher dependency between $Z_{i}$ and the knowledge selection behaviour, we sorted $\{I(Z_{i};Y)\}_{i=1}^{n}$ in descending order, and consider the top $k$ activation work as functional activations. To decide the number of selected activations $k$, we introduce a hyperparameter $K$. We then choose $k$ such that:
@@ C.3 Impact of the Size of the Collected Activations
(a) Use different numbers of activations to calculate mutual information. (Section 4.2)
(b) Use different numbers of activations to calculate $\mathbf{z}_{C}$ and $\mathbf{z}_{M}$. (Section 4.1)
@@ Appendix F Distribution Patterns of the Residual Stream Under Knowledge Conflict
In this section, we provide further analysis of the representation patterns when knowledge conflicts in addition to Fig. 5(d). Here, we focus on analysing the distribution patterns of different knowledge selection behaviours. More specifically, we compare the representation difference between the activation from $\mathcal{D}_{C}$ and $\mathcal{D}_{M}$, which are both under knowledge conflict but select contextual and parametric knowledge to generate the answer $C$ and $M$. In Section F.1, we analyse the skewness patterns; in Section F.2, we analyse the L1 norm and L2 norm patterns since previous work (Devoto et al., 2024) also show the norm value may be related to the contextual information usage. We provide more residual stream analysis in our preliminary study (Zhao et al., 2024b).